OpenStudy (anonymous):

The distance you travel while hiking is a function of how fast you hike and how long you hike at this rate. You usually maintain a speed of three miles per hour while hiking. Write a statement that describes how the distance that you travel is determined. Then identify the independent and dependent variables of this function. a.The distance traveled is three times the number of hours I have hiked. The independent variable is hours. The dependent variable is distance. b.The distance traveled is three times the number of hours I have hiked. The independent variable is distance. The dependent

OpenStudy (anonymous):

....The dependent variable is hours.. c. The hours I have hiked is three times the distance. The independent variable is distance. The dependent variable is hours. d. The hours I have hiked is three times the distance. The independent variable is hours. The dependent variable is distance

OpenStudy (anonymous):

Use the rate equation: \[rate=\frac{distance}{time}\] and solve for distance: \[distance=rate \times time\] and then substitute the known rate: \[distance = 3\frac{miles}{hour} \times time\] So right away the answer is A or B. The dependent variable is the input and the independent variable is the output so the answer is B. i.e. Distance depends on time.

OpenStudy (anonymous):

thank you so much @twvogels

OpenStudy (anonymous):

The distance you travel while hiking is a function of how fast you hike and how long you hike at this rate. You usually maintain a speed of three miles per hour while hiking. Write an equation that represents this function, letting t represent the indpendent variable, let h represent the function, and h(t) the dependent variable. Then, determine the h(4). a. h(t) = 3t/4 ; 48 miles b. t(h) = 3h ; 12 miles c. h(t) = 3t ; 12 miles d. h(t) = 3t/4 ; 3 miles

OpenStudy (anonymous):

can you help me with this one? @twvogels

OpenStudy (anonymous):

\[h(t)=3t\]\[h(4)=3(4)=12\] C

OpenStudy (anonymous):

ohh i understand how to do it now! thank you so much :) @twvogels

OpenStudy (anonymous):

The distance you travel while hiking is a function of how fast you hike and how long you hike at this rate. You usually maintain a speed of three miles per hour while hiking. Write a statement using function notation; let t represent the independent variable. Let h represent the function, and h(t) the dependent variable. a. h(t) = 3h b. h(t) = 3t c.t(h) = 3t d. t(h) = 3h please help

OpenStudy (anonymous):

Same as above \[h(t)=3t\]

OpenStudy (anonymous):

If you work for an hourly wage, your gross pay is a function of the number of hours that you work. If you earn $8 per hour, complete the following table: Number of Hours 0 3 5 7 10 12 Gross Pay If you were to plot the ordered pairs on a set of axes, in which quadrant would the domain and range of the function lie? a. 8, 16, 24, 32, 40, 48; Quadrant I b. 0, 24, 40, 56, 80, 96; Quadrant II c. 8, 16, 24, 32, 40, 48; Quadrant II d. 0, 24, 40, 56, 80, 96; Quadrant I

OpenStudy (anonymous):

the number of hours and gross pay is in a table. Number Of Hours 0 3 5 7 10 12 Gross Pay

OpenStudy (anonymous):

Number of Hours Gross Pay 0 0 3 24 5 40 7 56 10 80 12 96 All the values a zero or positive, so the function lies in Quad 1. D

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You're welcome.